Can you solve a system of equations of
4=a(b)^0
1=a(b)^2

Answer by mananth(16949)

4=a(b)^0
1=a(b)^2

b^0=1
therefore a=4
1=a(b)^2
1=4(b)^2
1/4 = b^2
b= +/-(1/2)